# CRISP-SAM2 : SAM2 with Cross-Modal Interaction and Semantic Prompting for Multi-Organ Segmentation

### v2

@@ 1. Introduction
While SAM, SAM2, and their variants have achieved significant success, there remains room for improvement, particularly in medical multi-organ segmentation. One of the primary challenges is related to their pre-training process (Xin et al., 2024c). Although these models are trained on extensive datasets featuring natural objects, the scarcity of medical data within these sets hinders the ability of the model to learn specific features of medical imaging effectively. As a result, the gap between the natural and medical domains remains unbridged. To address this, some researchers (Ma et al., 2024; Wang et al., 2025) have adapted SAM for medical segmentation by training it on large-scale medical datasets. Other models (Wei et al., 2024; Guo et al., 2025; Zhu et al., 2024; Huang et al., 2024a) have modified the original structures or added additional components to better accommodate the unique characteristics of medical imaging, thereby improving performance. Recently, innovative models (Huang et al., 2025; Jiang et al., 2024; Du et al., 2025) have begun to explore the integration of textual information to assist in segmentation. The rich contextual information inherent in natural language offers great potential for enhancing segmentation performance, pointing to promising directions for future research.
@@
To address the limitations mentioned earlier, we first introduce a two-level progressive cross-attention structure to generate cross-modal semantics. These semantics are integrated into the image encoder as contextualized information to enhance the understanding of visual features in detail. By converting these cross-modal semantics and image features into sparse and dense prompt embeddings, we can eliminate the dependency on geometric prompts. Besides, this semantic prompting method, which is guided by textual descriptions of organs, allows for capturing localized details and complex boundaries with improved accuracy. In addition to the original mask, we reuse the mask decoder on an additional learnable tokens and generate a refined mask, which is combined with the cross-modal semantics to further refine segmentation details. Recognizing the unique characteristics of 3D medical imaging and the ability of SAM2 to segment continuous imaging frames, we replace its original first-in-first-out (FIFO) strategy with a similarity-sorting self-updating strategy to better leverage spatial information. Our proposed model, CRISP-SAM2, based on SAM as shown in Figure 1, outperforming both purely visual models (Ma et al., 2024; Wei et al., 2024; Zhu et al., 2024; Guo et al., 2025) and text-assisted models (Li et al., 2023a; Huang et al., 2025; Du et al., 2025; Jiang et al., 2024) on seven public datasets. Comprehensive ablation studies and visualizations confirm the effectiveness and advantages of our designs. The key contributions of this work include:
@@ 3.2. Semantic Extraction
Recognizing that are unseen to the model, we employ a Cross-Modal Semantic Injector, i.e. CS-Injector, to integrate this complementary feature into the pre-trained image encoder of SAM2 (Ma et al., 2024). This approach largely follows the architecture of SAM2-Adapter (Chen et al., 2023, 2024d). Specifically, we frozen the pre-trained weights of the Hiera image encoder to preserve its strong visual representations. In addition, a simple and efficient CS-Injector is designed to inject semantics, consisting of just two light-weight MLPs. The dimensions of the second MLP correspond to the four hierarchical transformer blocks within the image encoder. It is worth noting that we additionally sum the output of the first MLP with the original input of visual features as the input of the second MLP. Ultimately, the injected image encoder produces the Image Embedding .
@@ 4.2. Comparison Results
To evaluate the performance of our CRISP-SAM2 model, we perform comparative experiments against other SOTA methods in the seven datasets, utilizing SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2025) models as baselines. Our comparisons include both purely visual segmentation approaches, namely MedSAM (Ma et al., 2024), I-MedSAM (Wei et al., 2024), MedSAM-2 (Zhu et al., 2024), and CT-SAM3D (Guo et al., 2025), and text-engaged segmentation methods, including LViT (Li et al., 2023a), CAT (Huang et al., 2025), SegVol (Du et al., 2025), and ZePT (Jiang et al., 2024). For the comparison of prompt-required models, we utilize three points and a bounding box as prompts, which are randomly selected from the center of ground truth and a rectangle completely covers targets. More comparison results are shown in Appendix E.1.
@@
Compared to unimodal medical segmentation methods (Ma et al., 2024; Wei et al., 2024; Zhu et al., 2024; Guo et al., 2025), CRISP-SAM2 achieves the highest DSC and NSD values across all datasets, with at least 2.99% and 4.03% enhancement in DSC and NSD evenly. Among all the text-guided methods (Li et al., 2023a; Huang et al., 2025; Du et al., 2025; Jiang et al., 2024), CRISP-SAM2 demonstrates the best overall performance, with an average improvement of at least 0.74% in DSC and 1.62% in NSD across the seven datasets. Specifically, on the AbdomenCT-1k (Ma et al., 2021) dataset, which is the largest of the seven, with 800 training samples and 200 test samples, CRISP-SAM2 outperforms the four relatively outstanding comparative methods, i.e. CT-SAM3D (Guo et al., 2025), CAT (Huang et al., 2025), SegVol (Du et al., 2025) and ZePT (Jiang et al., 2024) with DSC and NSD improvements of 2.23%/3.90%/4.58%/0.94% and 1.85%/3.61%/5.05%/2.03%, respectively, demonstrating superior segmentation accuracy.
@@ A.1. SAM and SAM2
Deep learning has made impressive progress in several fields, such as computer vision (Lu et al., 2023; Li et al., 2025c, f, 2024a; Wu et al., 2025d; Song et al., 2025a; Zeng et al., 2024; Chen et al., 2024a; Lu et al., 2024a; Qian et al., 2024; Yao et al., 2025; Xiao et al., 2025a; Song et al., 2025b; Li et al., 2024b; Lu et al., 2024b; Yu et al., 2025; Li et al., 2024c; Gao et al., 2024; Zeng et al., 2025; Yao and Chuah, 2024; Li et al., 2025b; Song et al., 2024), natural language processing (Bi et al., 2024, 2025; Chen et al., 2024c; Li et al., 2024e; Xiao et al., 2025b), automated control (Zhou et al., 2025a, b), time series analysis (Qiu et al., 2024; Wu et al., 2025b; Qiu et al., 2025c; Wu et al., 2025c; Qiu et al., 2025b; Liu et al., 2025a; Qiu et al., 2025a), multi-modal analysis (Wu et al., 2025a; Gu et al., 2025; Li et al., 2025e; Wu et al., 2025e; Li et al., 2025d), and multi-task learning (Xin et al., 2024a, b). However, there is still a lack of unified models in the field of computer vision, which is the reason why SAM and SAM2 are proposed.
@@ C.2. Implementation Details
Following the data augmentation in previous studies (Zhao et al., 2024; Guo et al., 2025; Huang et al., 2025), random shift and rotation, gamma scaling, and brightness adjustment operations are applied to expand training samples and balance the numbers of each class of organ. We choose Hiera-B+ as the backbone and for the text encoder and image encoder in the Cross-Modal Semantic Interaction module, we adopt ViT-B/16 and CLIP-ViT-B/16 respectively, and fix their parameters. The input imaging is resized to for the image encoder in this module. As listed in Table 7 all experiments are conducted on 8 NVIDIA A100-SXM4-80G GPUs in the PyTorch (Paszke et al., 2019) framework with AdamW (Loshchilov and Hutter, 2017) optimizer. To make the best use of the global information of visual inputs under the limitations of the memory of GPUs, the input patch size and batch size are set to and 16, with 2 batches per GPU.
@@ E.1. Comparison Results
In addition to the experimental results elaborated in Section 4.2, we employ box plots comparing the DSC and NSD metrics on each dataset to provide complementary evidence about the performances of other SOTA methods and our CRISP-SAM2. As illustrated in Figure 7, our CRISP-SAM2 significantly outperforms other counterparts. In summary, the medians of our method, which is denoted by the dashed lines in each plot, exceed those of other methods in twelve of these fourteen charts. Besides, the size of the rectangle of the box plot embodies the degree of disaggregation of the prediction metrics and segmentation stability of the models. Our CRISP-SAM2 has the smallest rectangles in all plots, demonstrating stable performance on various samples and better robustness. Together with the average values listed in Table 1, the box plots verify our superior performance in terms of the metrics of segmentation effectiveness. To offer more intuitive evidence, we visualize more detailed predicted segmentation results of CT-SAM3D (Guo et al., 2025), CAT (Huang et al., 2025), SegVol (Du et al., 2025), ZePT (Jiang et al., 2024) and our CRISP-SAM2. And we provide the box plots in Figure 8 for the average values of each organ on the seven selected datasets, showing segmentation performance differences between organs because of their unique characteristics. As depicted in Figure 9, the detailed visualization zoom-in intuitively exhibited that our segmentation results are more closely aligned to the ground truths, especially the more accurate local details and boundaries.
@@ E.2. Zero-Shot Generalization Performance
Since the powerful zero-shot generalization ability of SAM2, which could be further enhanced by geometric prompts, we conduct experiments to compare its zero-shot generalization performance with our model. We choose colon, intestine, and rectum in WORD (Luo et al., 2022), which are not included in training, as ”Unseen” labels 1 to 3 respectively for evaluation. For models requiring geometric prompts, i.e. SAM2 (Ravi et al., 2025), MedSAM-2 (Zhu et al., 2024), CT-SAM3D (Guo et al., 2025) and SegVol (Du et al., 2025), we provide points and bboxs, following the previous experiments. As listed in Table 8, it could be observed that these geometric-prompted models have smaller performance reduction when segmenting unseen targets than those without geometric prompts, namely CAT (Huang et al., 2025), ZePT (Jiang et al., 2024) and our model. It is mainly because that geometric prompts are more direct and spatially precise compared to our text-based prompting strategy, requiring exploring inherent relationships between vision and text. While CRISP-SAM2 slightly falls behind SegVol by 0.72% on the average value of DSC, it has a lead of 0.22% on NSD evenly when facing the three unseen organs. Compared to baseline SAM2, our method surpasses it 3.53% and 7.70% on DSC and NSD respectively. These results demonstrate that our CRISP-SAM2, which utilizes cross-modal interaction and semantic prompting strategies, is capable of generalization performance.

### v3

@@ 1. Introduction
While SAM, SAM2, and their variants have achieved significant success, there remains room for improvement, particularly in medical multi-organ segmentation. One of the primary challenges is related to their pre-training process (Xin et al., 2024c). Although these models are trained on extensive datasets featuring natural objects, the scarcity of medical data within these sets hinders the ability of the model to learn specific features of medical imaging effectively. As a result, the gap between the natural and medical domains remains unbridged. To address this, some researchers (Ma et al., 2024; Wang et al., 2025) have adapted SAM for medical segmentation by training it on large-scale medical datasets. Other models (Wei et al., 2024; Guo et al., 2025b; Zhu et al., 2024; Huang et al., 2024a) have modified the original structures or added additional components to better accommodate the unique characteristics of medical imaging, thereby improving performance. Recently, innovative models (Huang et al., 2025; Jiang et al., 2024; Du et al., 2025) have begun to explore the integration of textual information to assist in segmentation. The rich contextual information inherent in natural language offers great potential for enhancing segmentation performance, pointing to promising directions for future research.
@@
To address the limitations mentioned earlier, we first introduce a two-level progressive cross-attention structure to generate cross-modal semantics. These semantics are integrated into the image encoder as contextualized information to enhance the understanding of visual features in detail. By converting these cross-modal semantics and image features into sparse and dense prompt embeddings, we can eliminate the dependency on geometric prompts. Besides, this semantic prompting method, which is guided by textual descriptions of organs, allows for capturing localized details and complex boundaries with improved accuracy. In addition to the original mask, we reuse the mask decoder on an additional learnable tokens and generate a refined mask, which is combined with the cross-modal semantics to further refine segmentation details. Recognizing the unique characteristics of 3D medical imaging and the ability of SAM2 to segment continuous imaging frames, we replace its original first-in-first-out (FIFO) strategy with a similarity-sorting self-updating strategy to better leverage spatial information. Our proposed model, CRISP-SAM2, based on SAM as shown in Figure 1, outperforming both purely visual models (Ma et al., 2024; Wei et al., 2024; Zhu et al., 2024; Guo et al., 2025b) and text-assisted models (Li et al., 2023a; Huang et al., 2025; Du et al., 2025; Jiang et al., 2024) on seven public datasets. Comprehensive ablation studies and visualizations confirm the effectiveness and advantages of our designs. The key contributions of this work include:
@@ 3.2. Semantic Extraction
Recognizing that are unseen to the model, we employ a Cross-Modal Semantic Injector, i.e. CS-Injector, to integrate this complementary feature into the pre-trained image encoder of SAM2 (Ma et al., 2024). This approach largely follows the architecture of SAM2-Adapter (Chen et al., 2023c, 2024d). Specifically, we frozen the pre-trained weights of the Hiera image encoder to preserve its strong visual representations. In addition, a simple and efficient CS-Injector is designed to inject semantics, consisting of just two light-weight MLPs. The dimensions of the second MLP correspond to the four hierarchical transformer blocks within the image encoder. It is worth noting that we additionally sum the output of the first MLP with the original input of visual features as the input of the second MLP. Ultimately, the injected image encoder produces the Image Embedding .
@@ 4.2. Comparison Results
To evaluate the performance of our CRISP-SAM2 model, we perform comparative experiments against other SOTA methods in the seven datasets, utilizing SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2025) models as baselines. Our comparisons include both purely visual segmentation approaches, namely MedSAM (Ma et al., 2024), I-MedSAM (Wei et al., 2024), MedSAM-2 (Zhu et al., 2024), and CT-SAM3D (Guo et al., 2025b), and text-engaged segmentation methods, including LViT (Li et al., 2023a), CAT (Huang et al., 2025), SegVol (Du et al., 2025), and ZePT (Jiang et al., 2024). For the comparison of prompt-required models, we utilize three points and a bounding box as prompts, which are randomly selected from the center of ground truth and a rectangle completely covers targets. More comparison results are shown in Appendix E.1.
@@
Compared to unimodal medical segmentation methods (Ma et al., 2024; Wei et al., 2024; Zhu et al., 2024; Guo et al., 2025b), CRISP-SAM2 achieves the highest DSC and NSD values across all datasets, with at least 2.99% and 4.03% enhancement in DSC and NSD evenly. Among all the text-guided methods (Li et al., 2023a; Huang et al., 2025; Du et al., 2025; Jiang et al., 2024), CRISP-SAM2 demonstrates the best overall performance, with an average improvement of at least 0.74% in DSC and 1.62% in NSD across the seven datasets. Specifically, on the AbdomenCT-1k (Ma et al., 2021) dataset, which is the largest of the seven, with 800 training samples and 200 test samples, CRISP-SAM2 outperforms the four relatively outstanding comparative methods, i.e. CT-SAM3D (Guo et al., 2025b), CAT (Huang et al., 2025), SegVol (Du et al., 2025) and ZePT (Jiang et al., 2024) with DSC and NSD improvements of 2.23%/3.90%/4.58%/0.94% and 1.85%/3.61%/5.05%/2.03%, respectively, demonstrating superior segmentation accuracy.
@@ A.1. SAM and SAM2
Deep learning has made impressive progress in several fields, such as computer vision (Lu et al., 2023; Li et al., 2025c, f; Chen and Li, 2025; Li et al., 2024a; Wu et al., 2025d; Song et al., 2025a; Zeng et al., 2024a; Zhang et al., 2025b; Chen et al., 2024a; Lu et al., 2024a; Qian et al., 2024; Yao et al., 2025; Chen et al., 2023b; Xiao et al., 2025a; Song et al., 2025b; Li et al., 2024b; Lu et al., 2024b; Yu et al., 2025; Li et al., 2024c; Gao et al., 2024; Zhang et al., 2025a; Cai et al., 2024; Chen et al., 2024e; Zeng et al., 2025; Yao and Chuah, 2024; Li et al., 2025b; Song et al., 2024), natural language processing (Bi et al., 2024, 2025; Chen et al., 2024c, 2023a; Li et al., 2024e; Xiao et al., 2025b), time series analysis (Qiu et al., 2024; Wu et al., 2025b; Qiu et al., 2025c; Wu et al., 2025c; Qiu et al., 2025b; Liu et al., 2025a; Qiu et al., 2025a), multi-modal analysis (Wu et al., 2025a; Zeng et al., 2024b; Guo et al., 2025a; Gu et al., 2025; Li et al., 2025e; Wu et al., 2025e; Li et al., 2025d), and multi-task learning (Xin et al., 2024a, b; Wu et al., 2024, 2025f). However, there is still a lack of unified models in the field of computer vision, which is the reason why SAM and SAM2 are proposed.
@@ C.2. Implementation Details
Following the data augmentation in previous studies (Zhao et al., 2024; Guo et al., 2025b; Huang et al., 2025), random shift and rotation, gamma scaling, and brightness adjustment operations are applied to expand training samples and balance the numbers of each class of organ. We choose Hiera-B+ as the backbone and for the text encoder and image encoder in the Cross-Modal Semantic Interaction module, we adopt ViT-B/16 and CLIP-ViT-B/16 respectively, and fix their parameters. The input imaging is resized to for the image encoder in this module. As listed in Table 7 all experiments are conducted on 8 NVIDIA A100-SXM4-80G GPUs in the PyTorch (Paszke et al., 2019) framework with AdamW (Loshchilov and Hutter, 2017) optimizer. To make the best use of the global information of visual inputs under the limitations of the memory of GPUs, the input patch size and batch size are set to and 16, with 2 batches per GPU.
@@ E.1. Comparison Results
In addition to the experimental results elaborated in Section 4.2, we employ box plots comparing the DSC and NSD metrics on each dataset to provide complementary evidence about the performances of other SOTA methods and our CRISP-SAM2. As illustrated in Figure 7, our CRISP-SAM2 significantly outperforms other counterparts. In summary, the medians of our method, which is denoted by the dashed lines in each plot, exceed those of other methods in twelve of these fourteen charts. Besides, the size of the rectangle of the box plot embodies the degree of disaggregation of the prediction metrics and segmentation stability of the models. Our CRISP-SAM2 has the smallest rectangles in all plots, demonstrating stable performance on various samples and better robustness. Together with the average values listed in Table 1, the box plots verify our superior performance in terms of the metrics of segmentation effectiveness. To offer more intuitive evidence, we visualize more detailed predicted segmentation results of CT-SAM3D (Guo et al., 2025b), CAT (Huang et al., 2025), SegVol (Du et al., 2025), ZePT (Jiang et al., 2024) and our CRISP-SAM2. And we provide the box plots in Figure 8 for the average values of each organ on the seven selected datasets, showing segmentation performance differences between organs because of their unique characteristics. As depicted in Figure 9, the detailed visualization zoom-in intuitively exhibited that our segmentation results are more closely aligned to the ground truths, especially the more accurate local details and boundaries.
@@ E.2. Zero-Shot Generalization Performance
Since the powerful zero-shot generalization ability of SAM2, which could be further enhanced by geometric prompts, we conduct experiments to compare its zero-shot generalization performance with our model. We choose colon, intestine, and rectum in WORD (Luo et al., 2022), which are not included in training, as ”Unseen” labels 1 to 3 respectively for evaluation. For models requiring geometric prompts, i.e. SAM2 (Ravi et al., 2025), MedSAM-2 (Zhu et al., 2024), CT-SAM3D (Guo et al., 2025b) and SegVol (Du et al., 2025), we provide points and bboxs, following the previous experiments. As listed in Table 8, it could be observed that these geometric-prompted models have smaller performance reduction when segmenting unseen targets than those without geometric prompts, namely CAT (Huang et al., 2025), ZePT (Jiang et al., 2024) and our model. It is mainly because that geometric prompts are more direct and spatially precise compared to our text-based prompting strategy, requiring exploring inherent relationships between vision and text. While CRISP-SAM2 slightly falls behind SegVol by 0.72% on the average value of DSC, it has a lead of 0.22% on NSD evenly when facing the three unseen organs. Compared to baseline SAM2, our method surpasses it 3.53% and 7.70% on DSC and NSD respectively. These results demonstrate that our CRISP-SAM2, which utilizes cross-modal interaction and semantic prompting strategies, is capable of generalization performance.